[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Mon Jul 18, 2016

HSU Shift Log nhc.tsbadmin at noaa.gov
Tue Jul 19 03:02:31 UTC 2016


HSU Shift Log for Mon Jul 18 2016 23:02:24 GMT-0400 (EDT)

Shift: R

Forecaster: E Blake

Technical Issues Summary:

ATCF issues: TAFB fix still not importing in a timely fashion. 00z fix came  
in at 01z. This is also important because for invest cases, if the fixes  
aren't in the decks at the time that the guidance is run, the TCGI guidance  
will fail.

N-AWIPS/AWIPS2 issues: None

Other Techincal issues: None
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20160719/56f9ea8e/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list